# Integral bases, perfect matchings, and the Petersen graph

### v2

@@ Appendix A Details about the composition procedure
Let us provide a proof of 4.1, parts (1)-(3). Our proof is very similar, and at times identical, to that of ([1], Lemma 5.1). First, it is clear from the construction that $B\subseteq P(G;C)\cap\{0,1\}^{E}$, and $|B|=|B_{1}|+|B_{2}|-|C|$. We need the following two claims.
@@
where (1) computes the coefficient for $x^{i_{1}}\odot y$, while (2) computes the coefficients for $x^{i_{t}}\odot y,t=2,\ldots,k$. Similarly, given that $B_{2}$ is linearly independent, for each $y^{j_{t}}$, the sum of the coefficients of vectors in $B$ of the form $x\odot y^{j_{t}}$ for some $x$, must be 0. Subsequently, we obtain that
$\sum_{i=\ell}^{\ell+k-1}\lambda_{i}^{e}$ $=0$ (3)
@@
where (3) computes the coefficient for $x\odot y^{j_{1}}$, while (4) computes the coefficients for $x\odot y^{j_{t}},t=2,\ldots,\ell$. Observe that (1) and (4) imply that $\lambda_{1}^{e}=0$, so together with (2), we obtain that $\lambda^{e}_{i}=0$ for all $1\leq i\leq k+\ell-1$. As this holds for all $e\in C$, $\lambda^{e}_{i}=0$ for all $e\in C,1\leq i\leq|I_{e}|+|J_{e}|-1$. $\Diamond$
@@
Since $x_{e}=y_{e}$, it can be readily checked that $\alpha(x^{i_{1}})-\sum_{t=2}^{\ell}\beta(y^{j_{t}})=\beta(y^{j_{1}})-\sum_{t=2}^{k}\alpha(x^{i_{t}})$, so
$\sum_{i=\ell}^{\ell+k-1}\lambda_{i}^{e}$ $=\beta(y^{j_{1}}).$ (8)
@@
Let $f\in P(G;C)\cap\{0,1\}^{E}$. Then $f=x\odot y$ where $x\in P(G_{1})\cap\{0,1\}^{E(G_{1})},y\in P(G_{2})\cap\{0,1\}^{E(G_{2})}$. As $B_{i}$ is a basis for $\operatorname{lin}(P(G_{i}))$ for $i\in[2]$, it follows that $x$ ($y$) is a linear combination of the vectors in $B_{1}$ ($B_{2}$), so by A.2, $f=x\odot y$ is a linear combination of the vectors in $B$. This, together with A.1, implies that $B$ is a basis for $\operatorname{lin}(P(G;C))$, thus finishing the proof for 4.1, part (1). Furthermore, if $B_{i}$ is a lattice basis for $L(G_{i})$ for $i\in[2]$, then $x$ ($y$) is an integer linear combination of the vectors in $B_{1}$ ($B_{2}$), so by A.2, $f=x\odot y$ is an integer linear combination of the vectors in $B$. In this case, $B$ is a lattice basis for the lattice generated by $P(G;C)\cap\{0,1\}^{E}$, in turn proving 4.1, part (2).
Finally, assume that $B_{i}$ is an integral basis for $\operatorname{lin}(P(G_{i}))$ for $i\in[2]$. Take $f\in\operatorname{lin}(B)\cap\mathbb{Z}^{E}$. Observe that $f=x\odot y$, where $x\in\operatorname{lin}(B_{1})\cap\mathbb{Z}^{E(G_{1})}$ and $y\in\operatorname{lin}(B_{2})\cap\mathbb{Z}^{E(G_{2})}$. As $B_{i}$ is an integral basis, $x$ ($y$) must be an integer linear combination of the vectors in $B_{1}$ ($B_{2}$), so by A.2, $f=x\odot y$ is an integer linear combination of the vectors in $B$. Thus, $B$ is an integral basis for its linear hull, in turn proving 4.1, part (3).∎

### v3

@@ Appendix A Details about the composition procedure
Let us provide a proof of 3.1, parts (1)-(3). Our proof is very similar, and at times identical, to that of ([1], Lemma 5.1). First, it is clear from the construction that $B\subseteq P(G;C)\cap\{0,1\}^{E}$, and $|B|=|B_{1}|+|B_{2}|-|C|$. We need the following two claims.
@@
where (1) computes the coefficient sum for $x^{i_{1}}\odot y$, while (2) computes the coefficient sums for $x^{i_{t}}\odot y,t=2,\ldots,k$. Similarly, given that $B_{2}$ is linearly independent, for each $y^{j_{t}}$, the sum of the coefficients of vectors in $B$ of the form $x\odot y^{j_{t}}$ for some $x$, must be 0. Subsequently, we obtain that
$\lambda_{1}^{e}+\sum_{i=\ell+1}^{\ell+k-1}\lambda_{i}^{e}$ $=0$ (3)
@@
where (3) computes the coefficient sum for $x\odot y^{j_{1}}$, while (4) computes the coefficient sums for $x\odot y^{j_{t}},t=2,\ldots,\ell$. Observe that (1) and (4) imply that $\lambda_{1}^{e}=0$, so together with (2), we obtain that $\lambda^{e}_{i}=0$ for all $1\leq i\leq k+\ell-1$. As this holds for all $e\in C$, $\lambda^{e}_{i}=0$ for all $e\in C,1\leq i\leq|I_{e}|+|J_{e}|-1$. $\Diamond$
@@
Since $x_{e}=y_{e}$, it can be readily checked that $\alpha(x^{i_{1}})-\sum_{t=2}^{\ell}\beta(y^{j_{t}})=\lambda_{1}^{e}=\beta(y^{j_{1}})-\sum_{t=2}^{k}\alpha(x^{i_{t}})$, so
$\lambda_{1}^{e}+\sum_{i=\ell+1}^{\ell+k-1}\lambda_{i}^{e}$ $=\beta(y^{j_{1}}).$ (8)
@@
Let $f\in P(G;C)\cap\{0,1\}^{E}$. Then $f=x\odot y$ where $x\in P(G_{1})\cap\{0,1\}^{E(G_{1})},y\in P(G_{2})\cap\{0,1\}^{E(G_{2})}$. As $B_{i}$ is a basis for $\operatorname{lin}(P(G_{i}))$ for $i\in[2]$, it follows that $x$ ($y$) is a linear combination of the vectors in $B_{1}$ ($B_{2}$), so by A.2, $f=x\odot y$ is a linear combination of the vectors in $B$. This, together with A.1, implies that $B$ is a basis for $\operatorname{lin}(P(G;C))$, thus finishing the proof for 3.1, part (1). Furthermore, if $B_{i}$ is a lattice basis for $L(G_{i})$ for $i\in[2]$, then $x$ ($y$) is an integer linear combination of the vectors in $B_{1}$ ($B_{2}$), so by A.2, $f=x\odot y$ is an integer linear combination of the vectors in $B$. In this case, $B$ is a lattice basis for the lattice generated by $P(G;C)\cap\{0,1\}^{E}$, in turn proving 3.1, part (2).
Finally, assume that $B_{i}$ is an integral basis for $\operatorname{lin}(P(G_{i}))$ for $i\in[2]$. Take $f\in\operatorname{lin}(B)\cap\mathbb{Z}^{E}$. Observe that $f=x\odot y$, where $x\in\operatorname{lin}(B_{1})\cap\mathbb{Z}^{E(G_{1})}$ and $y\in\operatorname{lin}(B_{2})\cap\mathbb{Z}^{E(G_{2})}$. As $B_{i}$ is an integral basis, $x$ ($y$) must be an integer linear combination of the vectors in $B_{1}$ ($B_{2}$), so by A.2, $f=x\odot y$ is an integer linear combination of the vectors in $B$. Thus, $B$ is an integral basis for its linear hull, in turn proving 3.1, part (3).∎